OpenStudy (anonymous):

Having trouble with indefinite integration by substitution. See problem below:

OpenStudy (anonymous):

\[\int\limits \sqrt{t^3-1} t^2 dt\] I understand that \[u=t^3-1\] and that\[du = 3t^2 dt\]

OpenStudy (anonymous):

Where I get confused is this. How am I to get to \[\frac{1}{3} \int\limits \sqrt{u} du\]

OpenStudy (anonymous):

Solve for dt and plug both equations back into the first equation. the t^2 will cancel out, and since 1/3 is a constant, it can be taken out of the integral.

OpenStudy (anonymous):

What steps am I supposed to take to solve for dt? (Sorry, it's probably blatantly obvious to you).

OpenStudy (anonymous):

if \[du = 3t ^{2} dt\] then divide both sides by 3t^2 to get: \[du/3t ^{2} = dt\]

OpenStudy (anonymous):

AAAH! Ok, I'm going to give that a try...

OpenStudy (anonymous):

\[du = 3t^{2}dt\] re-write this equation in terms of dt. Then plug in \[dt = du/(3t^{2})\] in your indefinite integral and voila you have an intergral in terms of du. That's what gets the equation you asked about \[1/3\int\limits_{?}^{?}\sqrt{u}du\]

OpenStudy (anonymous):

Thank you both, from that I get\[\frac{2}{9}(t^3-1)^{\frac{3}{2}}+C\]